# RepLoRA: Reparameterizing Low-Rank Adaptation via the Perspective of Mixture of Experts

### v1

@@ A.2.2 Global Part
Given that $\Theta$ is a compact set, there exists a mixing measure $G^{\prime}$ in $\bar{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that one of the $G^{\prime}_{n}$’s subsequences converges to $G^{\prime}$. Since $\mathcal{D}_{2}(G^{\prime}_{n},\bar{G}_{*})>\varepsilon^{\prime}$, we obtain that $\mathcal{D}_{2}(G^{\prime},\bar{G}_{*})>\varepsilon^{\prime}$. Applying the Fatou’s lemma, we have
@@
Given the least square estimator $\widetilde{G}_{n}$ in equation (16), the convergence rate of the regression function estimation $f_{\widetilde{G}_{n}}(\cdot)$ to the true regression function $f_{\widetilde{G}_{*}}(\cdot)$ under the $L^{2}(\mu)$ norm is given by:
@@
Given the rate of convergence of the regression function estimator $f_{\widetilde{G}_{n}}$ in Proposition A.2, our goal is to demonstrate the following inequality:
@@
The proof of that inequality also proceeds into local and global parts, similar to that of Theorem 4.2.
@@ A.3.1 Local Part
Since $\mathcal{D}_{3n}\to 0$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\to\exp(c^{*}_{j})$, ${\bm{B}}_{n,i}\to{\bm{B}}_{j}^{*}$, and ${\bm{A}}_{n,i}\to{\bm{A}}_{j}^{*}$ for any index $i$ in the Voronoi cell $\mathcal{V}_{j}$ and for any index $j\in[L]$. Throughout this proof, we assume without loss of generality that $M^{0}_{K,j}=I_{\bar{d}}$ with a note that our techniques can be extended to the general setting of that matrix. Now, the proof of the local part is divided into three steps as follows:
@@
$\widetilde{A}_{n,1}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})$
$+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\big{)}({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X}+L_{n,1,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}\Big{]}+\widetilde{R}_{n,1}(\mathbb{X}),$
@@
$\widetilde{B}_{n,2}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})$
$+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}+\mathbb{X}^{\top}L_{n,3,j}\mathbb{X}(\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j}))^{2}+L_{n,4,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+\mathbb{X}^{\top}L_{n,5,j}\mathbb{X}(\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X})^{2}$
$+L_{n,6,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\Big{]}f_{\widetilde{G}_{n}}(\mathbb{X})+\widetilde{R}_{n,4}(\mathbb{X}),$
@@
$Q_{n}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})$
$+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\big{)}({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X}+L_{n,1,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}$
@@
$+L_{n,6,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\big{)}$
×$\times({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X}+L_{n,1,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}+L_{n,4,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})$
$+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,6,j}+L_{n,7,j}^{\top}\mathbb{X}\Big{]}-\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})$
$+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\Big{]}f_{\widetilde{G}_{n}}(\mathbb{X})-\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})$
$+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}+\mathbb{X}^{\top}L_{n,3,j}\mathbb{X}(\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j}))^{2}+L_{n,4,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+\mathbb{X}^{\top}L_{n,5,j}\mathbb{X}(\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X})^{2}$
@@
$\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),$
$\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\sigma_{2}({\bm{B}}^{*}_{j}),\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}e_{u},\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\mathbb{X}^{(v)}(\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j}))^{2}\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),$
@@
We will proof that when $n\to\infty$, at least one of the coefficients of these functions does not go to 0. Assume by contrary that all these coefficients of these linear independent functions go to 0. From equation (36), we obtain that $\widetilde{M}_{n,j,\alpha_{1},\alpha_{2}}/\mathcal{D}_{3n}$, $\widetilde{M}_{n,j,\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}}/\mathcal{D}_{3n}$, and $\widetilde{N}_{n,j}/\mathcal{D}_{3n}$ go to 0 for all the coefficients $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}\in\mathbb{N}^{\bar{d}\times\bar{d}}$ satisfying that $1\leq|\alpha_{1}|+|\beta_{1}|+|\alpha_{2}|+|\beta_{2}|\leq 2$.
@@
Now, for any index $j\in[L]$ such that $|\mathcal{V}_{j}|=1$, the limits $\widetilde{M}_{n,j,e_{u},0_{d}}/\mathcal{D}_{3n}\to 0$ lead to
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\|\Delta{\bm{A}}_{n,ij}\|_{1}}{\mathcal{D}_{3n}}\to 0.$
Due to the equivalence between the $\ell_{1}$-norm and the $\ell_{2}$-norm, this result directly implies that
@@ A.3.2 Global Part
Given that the parameter space $\Theta$ is a compact set, there exists a mixing measure $\widetilde{G}^{\prime}$ in $\widetilde{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that one of the $\widetilde{G}^{\prime}_{n}$’s subsequences converges to $\widetilde{G}^{\prime}$. Since $\mathcal{D}_{3}(\widetilde{G}^{\prime}_{n},\widetilde{G}_{*})>\varepsilon^{\prime}$ for all $n\geq 1$, we obtain that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})\geq\varepsilon^{\prime}$. An application of the Fatou’s lemma leads to
[…]
$\geq\int\liminf_{n\to\infty}\left\|f_{\widetilde{G}^{\prime}_{n}}(\mathbb{X})-f_{\widetilde{G}_{*}}(\mathbb{X})\right\|^{2}d\mu(\mathbb{X}).$
The above inequality indicates that $f_{\widetilde{G}^{\prime}}=f_{\widetilde{G}_{*}}$ for almost surely $\mathbb{X}$. It follows from the identifiability property that $\widetilde{G}^{\prime}\equiv\widetilde{G}_{*}$. It follows that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})=0$, contradicting the fact that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
@@
Recall from the setting that the samples $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathbb{R}^{\bar{d}}\times\mathbb{R}^{\bar{d}}$ are i.i.d. from the following regression model:
@@
Here, $p(Y_{i}|f_{G}({\bm{X}}_{i}),\sigma^{2}I_{\bar{d}})$ stands for multivariate Gaussian distribution with mean $f_{G}({\bm{X}})$ and covariance matrix $\sigma^{2}I_{\bar{d}}$. An application of Theorem 7.4 from [49] leads to
@@
where the notation $h$ stands for the Hellinger distance. As the Hellinger distance between two multivariate Gaussian distributions has closed-form expression, direct calculation yields that
@@
The conclusion of the proposition is achieved.
Appendix B Related Works
Parameter-Efficient Fine-tuning (PEFT).
With the recent rise of large models, PEFT methods are growing in popularity for their ability to fine-tune large-scale models by training a relatively small number of parameters for adapting to specific downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as adapter-based methods, which introduce additional trainable parameters to the frozen backbone. For example, Series Adapter [11] proposes adding linear modules in sequence to the existing layer, or Parallel Adapter [9] proposes integrating these modules in parallel. The second category of PEFT methods is Prompt-based methods that add extra trainable soft tokens, referred to as prompts, to the input [24, 42, 51]. A weakness of these methods is that they increase inference latency compared to the original model.
Low-Rank Adaptation [12]. LoRA and its variants are among the third category of the PEFT method, which is well-known for its simplicity and for not adding extra inference burden. To fine-tune the linear layers of a large model, LoRA applies low-rank matrices to approximate the weight changes and then merge them to the pre-trained weights for inference. A recent variant of LoRA is DoRA [31], which proposes to decompose the weight change into a learnable magnitude and directional component. Another example is AdaLoRA [56], which parameterizes the incremental updates in the form of singular value decomposition and prunes less significant singular values for more efficient updates. Orthogonal Fine-tuning (OFT) [40] exploits the orthogonal factorization to fine-tune diffusion models. Recently, VeRA [18] significantly reduces the number of trainable parameters compared to LoRA by using learnable scaling vectors to adjust a shared pair of frozen random matrices across layers. Notably, our method, which will be proposed in the following few sections, also falls within this category, and we validate its efficacy alongside LoRA and its variants through theoretical analysis and comprehensive experimentation.
Mixture of Experts. Building on the foundational concept of mixture models [15, 17], prior works by [6, 44] established the MoE layer as a key component for efficiently scaling model capacity. MoE models have since gained widespread attention for their adaptability across various domains, including large language models [4, 58], computer vision [43, 38], and multi-task learning [32]. Recent studies have investigated the convergence rates for expert estimation in MoE models, focusing on different assumptions and configurations of gating and expert functions. [10], assuming data from an input-free gating Gaussian MoE, demonstrated that expert estimation rates for maximum likelihood estimation depend on the algebraic independence of the expert functions. Similarly, employing softmax gating, [37, 34] found that expert estimation rates are influenced by the solvability of polynomial systems arising from the interaction between gating and expert parameters. More recently, [35, 36] utilized least square estimation to propose an identifiable condition for expert functions, particularly for feedforward networks with nonlinear activations. They showed that estimation rates are significantly faster under these conditions than models using polynomial experts.
Appendix C Experimental Details
C.1 Hyperparameters

### v2

@@ A.2.2 Global Part
Given that $\Theta$ is a compact set, there exists a mixing measure $G^{\prime}$ in $\bar{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that one of the $G^{\prime}_{n}$’s subsequences converges to $G^{\prime}$. Since $\mathcal{D}_{2}(G^{\prime}_{n},\bar{G}_{*})>\varepsilon^{\prime}$, we obtain that $\mathcal{D}_{2}(G^{\prime},\bar{G}_{*})>\varepsilon^{\prime}$. An application of the Fatou’s lemma leads to
@@
Given the least square estimator $\widetilde{G}_{n}$ in equation (16), the convergence rate of the regression function estimation $f_{\widetilde{G}_{n}}(\cdot)$ to the true regression function $f_{\widetilde{G}_{*}}(\cdot)$ under the $L^{2}(\mu)$ norm is parametric on the sample size, that is,
@@
Given rate of $f_{\widetilde{G}_{n}}$ in Proposition A.2, our goal is to demonstrate the following inequality:
@@
We divide the proof of the above inequality into local and global parts. Before going into the proof details, let us introduce some essential assumptions on the activation function $\sigma$.
Assumptions. We impose the following assumptions on the activation functions $\sigma_{1}$ and $\sigma_{2}$:
(A.2) (Uniform Lipschitz) Let $F({\bm{X}};{\bm{B}},{\bm{A}}):=\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}})\sigma_{1}({\bm{A}}))\mathbb{X})({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}})\sigma_{1}({\bm{A}}))\mathbb{X}$. Then, for any $\tau\in\{1,2\}$, we have
$\sum_{|\alpha|=\tau}\Bigg{|}\Big{(}\frac{\partial^{|\alpha|}F}{\partial{\bm{A}}^{\alpha_{1}}\partial{\bm{B}}^{\alpha_{2}}}({\bm{X}};{\bm{B}},{\bm{A}})$ $-\frac{\partial^{|\alpha|}F}{\partial{\bm{A}}^{\alpha_{1}}\partial{\bm{B}}^{\alpha_{2}}}({\bm{X}};{\bm{B}}^{\prime},{\bm{A}}^{\prime})\Big{)}\gamma^{\alpha}\Bigg{|}\leq C\|({\bm{B}},{\bm{A}})-({\bm{B}}^{\prime},{\bm{A}}^{\prime})\|^{\zeta}\|\gamma\|^{\tau},$
for any vector $\gamma\in\mathbb{R}^{2dr}$ and for some positive constants $\zeta$ and $C$ which are independent of ${\bm{X}}$ and $({\bm{B}},{\bm{A}}),({\bm{B}}^{\prime},{\bm{A}}^{\prime})$. Here, $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}^{r\times d}\times\mathbb{N}^{d\times r}$.
are linearly independent for almost surely ${\bm{X}}$.
@@ A.3.1 Local Part
Since $\mathcal{D}_{3n}\to 0$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\to\exp(c^{*}_{j})$, ${\bm{B}}_{n,i}\to{\bm{B}}_{j}^{*}$, and ${\bm{A}}_{n,i}\to{\bm{A}}_{j}^{*}$ for any $i\in\mathcal{V}_{j}$ and $j\in[L]$. Throughout this proof, we assume without loss of generality that $M^{0}_{K,j}=I_{\bar{d}}$ with a note that our techniques can be extended to the general setting of that matrix. Now, the proof of the local part is divided into three steps as follows:
@@
$\widetilde{A}_{n,1}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\big{)}$
$({\bm{M}}^{0}_{V}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X}+L_{n,1,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}\Big{]}+\widetilde{R}_{n,1}(\mathbb{X}),$
@@
$\widetilde{B}_{n,2}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}$
$+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}+\mathbb{X}^{\top}L_{n,7,j}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\Big{]}f_{\widetilde{G}_{n}}(\mathbb{X})+\widetilde{R}_{n,4}(\mathbb{X}),$
@@
$Q_{n}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\big{)}({\bm{M}}^{0}_{V}$ (34)
$+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X}+L_{n,1,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}\big{(}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}$ (35)
$+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,2,j}+L_{n,4,j}^{\top}\mathbb{X}\sigma_{2}({\bm{B}}^{*}_{j})+\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}L_{n,6,j}+L_{n,7,j}^{\top}\mathbb{X}\Big{]}$
$-\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\Big{]}f_{\widetilde{G}_{n}}(\mathbb{X})$
$-\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+\sigma_{2}({\bm{B}}_{j}^{*})\sigma_{1}({\bm{A}}_{j}^{*}))\mathbb{X})\Big{[}L_{n,1,j}^{\top}\mathbb{X}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})+L_{n,2,j}^{\top}\mathbb{X}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}$ (36)
@@
$\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j})\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\sigma_{2}({\bm{B}}^{*}_{j})$
$\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\sigma_{1}({\bm{A}}^{*}_{j})\mathbb{X}e_{u},\quad\widetilde{U}(\mathbb{X};{\bm{B}}_{j}^{*}{\bm{A}}_{j}^{*})\mathbb{X}^{(u)}\mathbb{X}^{(v)}(\mathbb{X}^{\top}\sigma_{2}({\bm{B}}^{*}_{j}))^{2}\widetilde{V}(\mathbb{X};{\bm{B}}^{*}_{j},{\bm{A}}^{*}_{j}),$
@@
We demonstrate that at least one of the coefficients of these independent functions does not go to 0 as $n\to\infty$. Assume by contrary that all these coefficients of these linear independent functions go to 0. From equation (37), we obtain that $\widetilde{M}_{n,j,\alpha_{1},\alpha_{2}}/\mathcal{D}_{3n}$, $\widetilde{M}_{n,j,\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}}/\mathcal{D}_{3n}$, and $\widetilde{N}_{n,j}/\mathcal{D}_{3n}$ go to 0 for all the coefficients $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}\in\mathbb{N}^{\bar{d}\times\bar{d}}$ satisfying that $1\leq|\alpha_{1}|+|\beta_{1}|+|\alpha_{2}|+|\beta_{2}|\leq 2$.
@@
Now, for any index $j\in[L]$ such that $|\mathcal{V}_{j}|=1$, the limits $\widetilde{M}_{n,j,e_{u},0_{d}}/\mathcal{D}_{3n}\to 0$ lead to $\frac{\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\|\Delta{\bm{A}}_{n,ij}\|_{1}}{\mathcal{D}_{3n}}\to 0$ as $n\to\infty$. Due to the equivalence between the $\ell_{1}$-norm and the $\ell_{2}$-norm, this result directly implies that
@@ A.3.2 Global Part
Given that $\Theta$ is a compact set, there exists a mixing measure $\widetilde{G}^{\prime}$ in $\widetilde{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that one of the $\widetilde{G}^{\prime}_{n}$’s subsequences converges to $\widetilde{G}^{\prime}$. Since $\mathcal{D}_{3}(\widetilde{G}^{\prime}_{n},\widetilde{G}_{*})>\varepsilon^{\prime}$, we obtain that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})>\varepsilon^{\prime}$. An application of the Fatou’s lemma leads to
$0=\lim_{n\to\infty}\|f_{\widetilde{G}^{\prime}_{n}}-f_{\widetilde{G}_{*}}\|_{L^{2}(\mu)}\geq\int\liminf_{n\to\infty}\left\|f_{\widetilde{G}^{\prime}_{n}}(\mathbb{X})-f_{\widetilde{G}_{*}}(\mathbb{X})\right\|^{2}d\mu(\mathbb{X}).$
The above inequality indicates that $f_{\widetilde{G}^{\prime}}=f_{\widetilde{G}_{*}}$ for almost surely $\mathbb{X}$. From the identifiability property, we deduce that $\widetilde{G}^{\prime}\equiv\widetilde{G}_{*}$. It follows that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})=0$, contradicting the fact that $\mathcal{D}_{3}(\widetilde{G}^{\prime},\widetilde{G}_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
@@
Recall from the setting that $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathbb{R}^{\bar{d}}\times\mathbb{R}^{\bar{d}}$ are i.i.d. samples from the following regression model:
@@
Here, $p(Y_{i}|f_{G}({\bm{X}}_{i}),\sigma^{2}I_{\bar{d}})$ stands for multivariate Gaussian distribution with mean $f_{G}({\bm{X}})$ and covariance matrix $\sigma^{2}I_{\bar{d}}$. An application of Theorem 7.4 from (van de Geer, 2000) leads to
@@
where $h$ stands for the Hellinger distance. As the Hellinger distance between two multivariate Gaussian distributions has closed-form expression, direct calculation yields that
@@
The proof of the proposition is completed.
Appendix B Experimental Details
B.1 Hyperparameters